# Comparing ImageNet Pre-training with Digital Pathology Foundation Models for Whole Slide Image-Based Survival Analysis

### v3

@@ 4.1 Implementation Details
In terms of datasets, we use WSIs from the publicly available TCGA database. Specifically, we use the Bladder Urothelial Carcinoma (BLCA) and the Breast Invasive Carcinoma (BRCA) cohorts in our experiments. We select FFPE H&E diagnostic slides for each cohort and we follow the convention of using one WSI per patient by selecting slides having "DX1" in their TCGA identifier. This selection process yields $N=373$ and $N=1061$ WSIs for BLCA and BRCA respectively. Survival analysis datasets comprise both censored and uncensored observations. Uncensored observations correspond to patients with known survival times, while censored observations pertain to patients for whom survival time is not recorded, often because the patient was alive at the time of analysis.
Table 1 lists the dimension of the feature vector produced by each backbone. We use the Negative Log Likelihood (NLL) as a bag-level loss function to train the aforementioned MIL networks and the concordance index to assess their prognostic accuracy. We divide the TCGA cohorts into train and validation sets and we report the results after applying $K$-fold cross validation ($K=5$). We list the hyperparameters used in Appendix 0.A.
@@ 4.2 Results
We report the $5-$fold validation results by listing the concordance index and the standard deviation for every TCGA cohort in Table 2. We also list the concordance index values achieved by 2 model ensembles, ResNet50 with UNI and UNI with Hibou-Base, in Table 3.
Datasets
BLCA BRCA Average
MIL Method ResNet50
MeanMIL [16] 0.536 $\pm$ 0.072 0.567 $\pm$ 0.061 0.552 $\pm$ 0.067
MaxMIL [16] 0.543 $\pm$ 0.044 0.570 $\pm$ 0.049 0.557 $\pm$ 0.047
ABMIL [7] 0.570 $\pm$ 0.069 0.565 $\pm$ 0.040 0.568 $\pm$ 0.055
TransMIL [8] 0.528 $\pm$ 0.049 0.591 $\pm$ 0.053 0.5595 $\pm$ 0.0511
MIL Method    UNI
MeanMIL [16] 0.570 $\pm$ 0.05 0.597 $\pm$ 0.085 0.584 $\pm$ 0.068
MaxMIL [16] 0.530 $\pm$ 0.042 0.610 $\pm$ 0.092 0.570 $\pm$ 0.067
ABMIL [7] 0.601 $\pm$ 0.046 0.616 $\pm$ 0.066 0.609 $\pm$ 0.056
TransMIL [8] 0.565 $\pm$ 0.014 0.628 $\pm$ 0.053 0.597 $\pm$ 0.038
MIL Method    Hibou-Base
MeanMIL [16] 0.572 $\pm$ 0.036 0.575 $\pm$ 0.089 0.574 $\pm$ 0.063
MaxMIL [16] 0.518 $\pm$ 0.031 0.550 $\pm$ 0.080 0.534 $\pm$ 0.056
ABMIL [7] 0.570 $\pm$ 0.043 0.617 $\pm$ 0.085 0.594 $\pm$ 0.064
TransMIL [8] 0.542 $\pm$ 0.053 0.557 $\pm$ 0.045 0.550 $\pm$ 0.049
@@
Datasets
BLCA BRCA Average
MIL Method ResNet50 + UNI
MeanMIL [16] 0.591 $\pm$ 0.028 0.581 $\pm$ 0.080 0.586 $\pm$ 0.060
MaxMIL [16] 0.527 $\pm$ 0.050 0.585 $\pm$ 0.089 0.556 $\pm$ 0.072
ABMIL [7] 0.596 $\pm$ 0.034 0.606 $\pm$ 0.063 0.601 $\pm$ 0.057
TransMIL [8] 0.58 $\pm$ 0.033 0.60 $\pm$ 0.078 0.593 $\pm$ 0.084
MIL Method    UNI + Hibou-Base
MeanMIL [16] 0.615 $\pm$ 0.020 0.634 $\pm$ 0.068 0.625 $\pm$ 0.050
MaxMIL [16] 0.544 $\pm$ 0.050 0.616 $\pm$ 0.069 0.580 $\pm$ 0.060
ABMIL [7] 0.587 $\pm$ 0.026 0.610 $\pm$ 0.008 0.599 $\pm$ 0.019
TransMIL [8] 0.551 $\pm$ 0.025 0.608 $\pm$ 0.058 0.580 $\pm$ 0.045
@@ 5 Conclusions
The results in Table 2 indicate that both histopathological feature extractors can consistently enhance the predictive prowess of the MIL networks used in this study. In most cases, the UNI backbone achieves a larger improvement over ResNet50 compared to Hibou-Base. The concordance index values reported in Table 3 reveal that an ensemble of the two digital pathology backbones consistently outperforms an ensemble consisting of a feature extractor pre-trained on ImageNet with one pre-trained on WSIs. Furthermore, it is worth noting the performance improvement of utilizing an ensemble of feature extractors. This enhancement is especially evident with MeanMIL, though the benefit diminishes with more complex MIL network architectures. It is also noteworthy that this improvement occurs despite the ensemble of ResNet50 and UNI having a larger embedding dimension than the UNI and Hibou-Base ensemble. Future work includes incorporating feature extractors pre-trained using different SSL approaches, besides self-distillation employed in DINOv2, to further assess the effect of different strategies on prognostic accuracy.
@@ 6 Acknowledgments
Appendix 0.A Hyperparameter Options
Datasets BLCA BRCA
Number of WSIs $373$ $1061$
Batch size $1$ $1$
Gradient accumulation steps $32$ $32$
Optimizer Adam Adam
Bag weight $0.7$ $0.7$
Epochs $200$ $200$
Early stopping Enabled Enabled
Earliest stop epoch $40$ $40$
L1 regularization $41e-$ 4 $41e-$ 4
Patience epochs $10$ $10$
Learning rate $41e-$ 4 $55e-$ 5
Table 4: Hyperparameter options common across all model configurations

### v4

@@ 4 Experiments & Results
4.1 Datasets
[…]
Figure 3: Distribution of censored and uncensored patients in each TCGA dataset
@@ 4.2 Implementation Details
Table 1 lists the dimension of the feature vector produced by each backbone along with the total number of parameters used during the pre-training phase. The same information for all model ensembles is provided in Table 2. We use the Negative Log Likelihood (NLL) as a bag-level loss function to train the aforementioned MIL networks and the concordance index to assess their prognostic accuracy. The number of trainable parameters for every MIL method is provided in Table A.1 in Appendix Appendix A: MIL Model Parameters. We divide the TCGA cohorts into train and validation sets and we report the results after applying $K$-fold cross validation ($K=5$).
Our codebase is inspired by the MambaMIL framework [9], and we build upon its publicly available implementation. The project is implemented in PyTorch, and our code repository can be found at https://github.com/MarioPaps/WSI-Survival-Analysis-ImageNet-vs-Digital-Pathology-Backbones.git.
For training, we use the Adam optimizer and we set the batch size equal to 111 because the number of patches extracted for each WSI varies. We therefore perform backpropagation using gradient accumulation with the number of steps set to 323232. Furthermore, we leverage early stopping and regularization to combat overfitting.
@@
Ensemble Combination Embedding Dimension
ResNet50 + UNI 2048
Hibou-Base + ResNet50 1792
Hibou-Base + UNI 1792
ResNet50 + UNI + Hibou-Base 2816
Table 2: Embedding Dimensions of the ensembles
@@ 4.3 Results
We report the 55-5 -fold validation results by listing the concordance index and the standard deviation for every TCGA cohort in Table 3. We also list the concordance index values achieved by 2 model ensembles, ResNet50 with UNI and UNI with Hibou-Base, in Table 4.
MIL Method BLCA LUAD BRCA Average
ResNet50
MeanMIL [16] 0.536 $\pm$ 0.072 0.580 $\pm$ 0.021 0.567 $\pm$ 0.061 0.561 $\pm$ 0.063
MaxMIL [16] 0.543 $\pm$ 0.044 0.596 $\pm$ 0.044 0.570 $\pm$ 0.049 0.570 $\pm$ 0.056
ABMIL [7] 0.570 $\pm$ 0.069 0.594 $\pm$ 0.031 0.565 $\pm$ 0.040 0.576 $\pm$ 0.061
TransMIL [8] 0.528 $\pm$ 0.049 0.597 $\pm$ 0.063 0.591 $\pm$ 0.053 0.560 $\pm$ 0.051
UNI
MeanMIL [16] 0.570 $\pm$ 0.05 0.603 $\pm$ 0.030 0.597 $\pm$ 0.085 0.590 $\pm$ 0.073
MaxMIL [16] 0.530 $\pm$ 0.042 0.625 $\pm$ 0.076 0.610 $\pm$ 0.092 0.588 $\pm$ 0.089
ABMIL [7] 0.601 $\pm$ 0.046 0.605 $\pm$ 0.061 0.616 $\pm$ 0.066 0.607 $\pm$ 0.071
TransMIL [8] 0.565 $\pm$ 0.014 0.602 $\pm$ 0.047 0.628 $\pm$ 0.053 0.598 $\pm$ 0.051
Hibou-Base
MeanMIL [16] 0.572 $\pm$ 0.036 0.614 $\pm$ 0.073 0.575 $\pm$ 0.089 0.587 $\pm$ 0.085
MaxMIL [16] 0.518 $\pm$ 0.031 0.605 $\pm$ 0.073 0.550 $\pm$ 0.080 0.558 $\pm$ 0.080
ABMIL [7] 0.570 $\pm$ 0.043 0.565 $\pm$ 0.080 0.617 $\pm$ 0.085 0.584 $\pm$ 0.088
TransMIL [8] 0.542 $\pm$ 0.053 0.611 $\pm$ 0.034 0.557 $\pm$ 0.045 0.570 $\pm$ 0.055
@@
MIL Method BLCA LUAD BRCA Average
ResNet50+UNI
MeanMIL [16] 0.591 $\pm$ 0.028 0.586 $\pm$ 0.064 0.581 $\pm$ 0.080 0.586 $\pm$ 0.075
MaxMIL [16] 0.527 $\pm$ 0.050 0.603 $\pm$ 0.049 0.585 $\pm$ 0.089 0.572 $\pm$ 0.080
ABMIL [7] 0.596 $\pm$ 0.034 0.600 $\pm$ 0.083 0.606 $\pm$ 0.063 0.601 $\pm$ 0.078
TransMIL [8] 0.580 $\pm$ 0.033 0.589 $\pm$ 0.050 0.600 $\pm$ 0.078 0.590 $\pm$ 0.070
UNI+Hibou-Base
MeanMIL [16] 0.615 $\pm$ 0.020 0.601 $\pm$ 0.059 0.634 $\pm$ 0.068 0.617 $\pm$ 0.065
MaxMIL [16] 0.544 $\pm$ 0.050 0.615 $\pm$ 0.032 0.616 $\pm$ 0.069 0.592 $\pm$ 0.064
ABMIL [7] 0.587 $\pm$ 0.026 0.623 $\pm$ 0.063 0.610 $\pm$ 0.008 0.607 $\pm$ 0.049
TransMIL [8] 0.551 $\pm$ 0.025 0.578 $\pm$ 0.060 0.608 $\pm$ 0.058 0.579 $\pm$ 0.062
ResNet50+Hibou-Base
MeanMIL [16] 0.603 $\pm$ 0.020 0.605 $\pm$ 0.058 0.656 $\pm$ 0.080 0.621 $\pm$ 0.071
MaxMIL [16] 0.530 $\pm$ 0.034 0.592 $\pm$ 0.046 0.563 $\pm$ 0.064 0.562 $\pm$ 0.061
ABMIL [7] 0.579 $\pm$ 0.034 0.604 $\pm$ 0.089 0.655 $\pm$ 0.078 0.613 $\pm$ 0.087
TransMIL [8] 0.561 $\pm$ 0.021 0.599 $\pm$ 0.050 0.643 $\pm$ 0.057 0.601 $\pm$ 0.056
ResNet50 + UNI + Hibou-Base
MeanMIL [16] 0.603 $\pm$ 0.022 0.595 $\pm$ 0.079 0.650 $\pm$ 0.074 0.616 $\pm$ 0.078
MaxMIL [16] 0.525 $\pm$ 0.044 0.603 $\pm$ 0.049 0.570 $\pm$ 0.063 0.566 $\pm$ 0.064
ABMIL [7] 0.598 $\pm$ 0.022 0.604 $\pm$ 0.096 0.644 $\pm$ 0.058 0.615 $\pm$ 0.081
TransMIL [8] 0.563 $\pm$ 0.036 0.590 $\pm$ 0.040 0.609 $\pm$ 0.053 0.616 $\pm$ 0.078
@@ 5 Conclusions
The results in Table 3 indicate that both histopathological feature extractors can consistently enhance the predictive prowess of the MIL networks used in this study. In most cases, the UNI backbone achieves a larger improvement over ResNet50 compared to Hibou-Base.
The reported concordance index values in Table 4 demonstrate that a significant performance improvement can be obtained by utilizing an ensemble of feature extractors. This improvement is particularly pronounced with the two simpler MIL architectures, which are MeanMIL [16] and MaxMIL [16]. Despite the ensemble of ResNet50 and UNI having a larger embedding dimension, the combinations of Hibou-Base with ResNet50 and UNI deliver more frequent improvements in outcome prediction accuracy. Furthermore, while combining two histopathological feature extractors typically results in a higher concordance index, mixing a feature extractor pre-trained on ImageNet with one pre-trained on WSIs outperforms the dual histopathological extractor approach in some cases. Lastly, the ensemble of all three feature extractors does not appear to yield a notable performance improvement in any MIL architecture.
Future work includes incorporating feature extractors pre-trained using different SSL approaches, besides self-distillation employed in DINOv2, to further assess the effect of different strategies on prognostic accuracy.
@@ 6 Acknowledgments
Appendix A: MIL Model Parameters
Table A.1 presents the number of trainable parameters for MeanMIL [16], MaxMIL [16], ABMIL [7], and TransMIL [8], assuming an input embedding dimension of 102410241024.
MIL Network Number of Parameters
MeanMIL [16] $526,852\approx 527K$
MaxMIL [16] $526,852\approx 527K$
ABMIL [7] $592,645\approx 593K$
TransMIL [8] $2,673,172\approx 2.7M$
Table A.1: Trainable parameter counts for MIL networks with an input embedding dimension of 1024.
Appendix B: Hyperparameter Choices
Table B.1 summarizes the hyperparameter choices that are consistently applied across all model setups.
Table B.1: Hyperparameter options common across all model configurations
Datasets BLCA LUAD BRCA
Number of WSIs 373373373 443443443 106110611061
Batch size 111 111 111
Gradient accumulation steps 323232 323232 323232
Optimizer Adam Adam Adam
Bag weight 0.70.70.7 0.70.70.7 0.70.70.7
[…]
Epochs 200200200 200200200 200200200
Early stopping Enabled Enabled Enabled
Earliest stop epoch 404040 404040 404040
[…]
Patience epochs 101010 555 101010
Dropout rate 0.25 0.25 0.25
[…]